OpenStudy (anonymous):

Using point-slope form, write the equation of the line that passes through the point (-4,12) and has a slope of -3/4.

OpenStudy (anonymous):

point slope formula is y-y1=m(x-x1). where m is the slope. in your case m=-3/4. replace the value of y1 and x1 for the corresponding point values, then solve for y.

OpenStudy (anonymous):

what...

OpenStudy (anonymous):

do you know the point-slope formula?

OpenStudy (anonymous):

not at all.

OpenStudy (anonymous):

never mind thats mx+b

OpenStudy (anonymous):

Okay, well the formula is the Following. \[y-y_{1}=m(x-x_{1})\]

OpenStudy (anonymous):

y=mx+b is called slope-intercept form. they're similar.

OpenStudy (anonymous):

okay. but i dont understand how to calculate it.

OpenStudy (anonymous):

Okay, well they give you a point (x,y). Which is (-4,12). so x = -4, y =12

OpenStudy (anonymous):

i know my points i just dont understand the question what so ever.

OpenStudy (anonymous):

so replace those two into the formula i gave you. and the m = -3/4

OpenStudy (anonymous):

so B is the answer

OpenStudy (anonymous):

No Ds the answer is im doing it correctly

OpenStudy (anonymous):

No, if you replace it you get (y-12)=-3/4(x+4)

OpenStudy (anonymous):

inside the parethesis you get (x-(-4)). Which is (x+4)

OpenStudy (anonymous):

alright thanks!

OpenStudy (anonymous):

No problem. Gluck!